Max wants a used car. He finds a deal for a Maxima at a total of $4200 if he can put 8.5% down and his equal monthly payments would be complete in one year. Max thinks it over and pays the down payment. How much are his monthly payments?
Answer by TimothyLamb(4379)

down:
4200 * 0.085 = 357.00
---
monthly payments:
4200 * (1 - 0.085) / 12 = 320.25
---
check:
320.25 * 12 + 357.00 = 4200
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php